OpenStudy (hungergames13):

Which ordered pair is the solution to the system of equations? x+y=4 x-y=-6 A. (2, 2) B. (1, −7) C. (−1, 5) D. (−2, 4)

OpenStudy (trojanpoem):

sum both equations: 2x = -2 x = -1 x - y = -6 -1 -y = -6 y = 5 solution set as (x, y) = (-1 , 5)

OpenStudy (trojanpoem):

You could also just plug in each set and check which will verify both of the equations , but that's just waste of time. x + 4 = -1 + 5 = 4 x - y = -1 -5 = -6